# Analytical results for the distribution of first return times of non-backtracking random walks on configuration model networks

### v1

@@ Abstract
We present analytical results for the distribution of first return (FR) times of non-backtracking random walks (NBWs) on undirected configuration model networks consisting of $N$ nodes with degree distribution $P(k)$. We focus on the case in which the network consists of a single connected component. Starting from a random initial node $i$ at time $t=0$, an NBW hops into a random neighbor of $i$ at time $t=1$ and at each subsequent step it continues to hop into a random neighbor of its current node, excluding the previous node. We calculate the tail distribution $P(T_{\rm FR}>t)$ of first return times from a random node to itself. It is found that $P(T_{\rm FR}>t)$ is given by a discrete Laplace transform of the degree distribution $P(k)$. This result exemplifies the relation between structural properties of a network, captured by the degree distribution, and properties of dynamical processes taking place on the network. Using the tail-sum formula, we calculate the mean first return time ${\mathbb{E}}[T_{\rm FR}]$. Surprisingly, ${\mathbb{E}}[T_{\rm FR}]$ coincides with the result obtained from the Kac’s lemma that applies to classical random walks (RWs). We also calculate the variance ${\rm Var}(T_{\rm FR})$, which accounts for the variability of first return times between different NBW trajectories. We apply this formalism to random regular graphs, Erdős-Rényi networks and configuration model networks with exponential and power-law degree distributions and obtain closed-form expressions for $P(T_{\rm FR}>t)$ as well as its mean and variance. These results provide useful insight on the advantages of NBWs over classical RWs in network exploration, sampling and search processes.
@@ 1 Introduction
Random walk (RW) models [1] provide useful tools for the analysis of dynamical processes on random networks [2, 3, 5, 4]. Here we focus on the case of undirected networks. Starting at time $t=0$ from a random initial node $i$, at each time step $t\geq 1$ an RW hops randomly to one of the neighbors of its current node. In some of the time steps the RW visits nodes that have not been visited before, while in other time steps it revisits nodes that have already been visited at an earlier time. The mean number $\langle S\rangle_{t}$ of distinct nodes visited by an RW on a random network up to time $t$ was recently studied [6]. It was found that in the infinite network limit, in which random networks exhibit a tree structure, $\langle S\rangle_{t}\simeq rt$, where the coefficient $r<1$ depends on the network topology. In this case, the revisits are due to backtracking steps in which the RW hops back to the previous node and subsequent retroceding steps in which it keeps hopping backwards along its own path [7].
@@
In order to perform systematic studies of random walks on random networks, it is useful to focus on configuration model networks. The configuration model is an ensemble of uncorrelated random networks consisting of $N$ nodes, whose degree sequences are drawn from a given degree distribution $P(k)$. The admissible degrees are often restricted to a finite range $k_{\rm min}\leq k\leq k_{\rm max}$, where $k_{\rm min}$ is the minimal degree and $k_{\rm max}$ is the maximal degree, such that for any value of $k$ outside this range $P(k)=0$. The mean degree $\langle K\rangle$ is denoted by $c$. To ensure that an RW starting from any initial node $i$ will be able to reach any other node $j$, we focus on the case in which the whole network consists of a single connected component. Using the terminology of percolation theory, these are networks in which the giant component encompasses the whole network. A sufficient condition for a configuration model network to consist of a single connected component is that $k_{\rm min}\geq 3$ [8]. In fact, single-component random networks also exist in the presence of nodes of degree $k=2$, as long as they are not too abundant. We thus avoid isolated nodes of degree $k=0$ and leaf nodes of degree $k=1$, which may form isolated tree structures.
@@
The first return (FR) time $T_{\rm FR}$ of an RW is the first time at which it returns to the initial node $i$ [9]. The first return time varies between different instances of the random walk trajectory and its properties can be captured by a suitable distribution. The distribution of first return times may depend on the specific realization of the random network and on the choice of the initial node $i$. The distribution of first return times from a random node to itself in a given ensemble of random networks is denoted by $P(T_{\rm FR}=t)$. A classical result regarding first return times is the Kac’s lemma, which states that the mean first return time of an RW from a given node $i$ to itself is given by [10, 11, 12]
@@
where $\langle X\rangle$ is the average of $X$ over the degree distribution $P(k)$.
@@
An important variant of the RW model is the non-backtracking random walk (NBW), in which the move backwards to the previous node is excluded [24]. Since backtracking steps are excluded, in the infinite network limit in which the network exhibits a tree structure, an NBW never revisits a previously visited node and $\langle S\rangle_{t}=t+1$. In particular, it never returns to the initial node. In a finite network, the first return process of NBWs takes place only via non-retroceding trajectories, which rely on the existence of cycles. Note that in networks that include leaf nodes of degree $k=1$, NBWs reaching these nodes would get stuck. Therefore, in the study of NBWs it is important not only to ensure that the network consists of a single connected component, but also that this component does not include any leaf nodes. This implies that the 2-core of the network (namely the largest subgraph in which every node has a degree of at least $2$) encompasses the whole network.
@@
NBWs are important for several reasons, which are summarized below. They provide a more efficient way to explore and analyze complex networks, compared to standard random walks. This is due to the fact that by avoiding backtracking steps they can cover more of the network in less time [25]. NBWs are useful for identifying community structures within networks [26]. In certain types of networks, NBWs can help mitigate localization effects that might trap standard random walks in specific regions of the network.
@@
NBWs belong to the class of second-order random walks, in which the transition probabilities depend not only on the current node but also on the previous node [27]. This introduces memory into the process, which makes it no longer Markovian in the traditional sense. The challenge is to analyze such processes using methods that are typically applied to Markov chains, which rely on the memoryless property. Recently, Fasino et al. introduced a mapping of second order random walks into first order processes on a larger state space, referred to as the pullback process [27]. Instead of viewing the random walk as taking place between the nodes of the original graph, the pullback process considers a random walk on the directed-line graph associated with the original graph. Using this method they showed that the mean first return time ${\mathbb{E}}[T_{\rm FR}]$ of any second order random walk (including NBWs) on undirected networks satisfies Eq. (1), thus extending the validity of the Kac’s lemma to second order random walks on undirected networks [27]. Note that the Kac’s lemma deals with the mean first return time and has no implications on the overall shape of the distribution and its higher order moments.
@@
NBWs exhibit faster mixing times than standard random walks, meaning they converge to their stationary distribution more quickly [28]. They thus inspire the design of more efficient algorithms for various graph-based problems, including link prediction and node centrality measures [29]. The non-backtracking (Hashimoto) matrix $B$ associated with these walks has spectral properties that can reveal important information about the network structure, often more clearly than traditional adjacency matrices [31, 30, 12]. Actually, the mixing time is inversely proportional to the spectral gap of the matrix $B$ [11]
@@
The paper is organized as follows. In Sec. 2 we derive formulae for the distribution $P(T_{\rm FR}>t)$ of first return times of NBWs on configuration model networks and for its mean ${\mathbb{E}}[T_{\rm FR}]$ and variance ${\rm Var}(T_{\rm FR})$. In Sec. 3 we apply these results to RRGs, Erdős-Rényi (ER) networks and to configuration model networks with exponential and power-law distributions. The results are discussed in Sec. 4 and summarized in Sec. 5.
@@ 2 The distribution of first return times
In the first three time steps the NBW crosses new edges, which implies that $\Delta L_{t}=1$ for $t=0,1$ and $2$. For $t\geq 3$ the probability $\Delta L_{t}$ is proportional to the fraction of yet uncrossed ’directed’ edges up to time $t$, namely
@@
Inserting $\Delta L_{t}$ from Eq. (5) into Eq. (6), we obtain a recursion equation for $\langle L\rangle_{t}$. Solving this equation, we obtain
@@
While Eq. (6) is valid to a good approximation for $t\geq 3$, it becomes precise above the mixing time, where the random walker samples edges in a uniform fashion. Thus, apart from the first few steps, Eq. (7) can be approximated by
@@
The probability that an NBW will not visit a specific random node of degree $k$ for the first time up to time $t$, can be expressed by $\left(1-\frac{\langle L\rangle_{t-1}}{Nc}\right)^{k}$. This is due to the fact that in order to visit a node of degree $k$ the NBW must enter via one of the $k$ edges connected to $i$. Since an NBW quickly loses memory of its initial node, the probability of not returning to an initial node of degree $k$ up to time $t$ is the same as the probability not to visit any other node of the same degree up to time $t$. Therefore, the tail distribution of first return times, under the condition that the initial node $i$ is of degree $k$, is given by
@@
and the second moment if given by
@@
This result coincides with the Kac’s lemma, which is obtained from general properties of discrete stochastic processes [10, 27]. Eq. (20) also implies that conditioning on initial nodes of a given degree $k$, the mean first return time is given by
@@
This results goes beyond the generalization of the Kac’s lemma for second-order random walks [27] and is valid for the specific case of the NBW.
@@
${\rm Var}(T_{\rm FR})=N^{2}c^{2}\left(2\bigg{\langle}\frac{1}{K^{2}}\bigg{\rangle}-\bigg{\langle}\frac{1}{K}\bigg{\rangle}^{2}\right).$ (25)

### v2

@@ Abstract
We present analytical results for the distribution of first return (FR) times of non-backtracking random walks (NBWs) on undirected configuration model networks consisting of $N$ nodes with degree distribution $P(k)$. We focus on the case in which the network consists of a single connected component. Starting from a random initial node $i$ at time $t=0$, an NBW hops into a random neighbor of $i$ at time $t=1$ and at each subsequent step it continues to hop into a random neighbor of its current node, excluding the previous node. We calculate the tail distribution $P(T_{\rm FR}>t)$ of first return times from a random initial node to itself. It is found that $P(T_{\rm FR}>t)$ is given by a discrete Laplace transform of the degree distribution $P(k)$. This result exemplifies the relation between structural properties of a network, captured by the degree distribution, and properties of dynamical processes taking place on the network. Using the tail-sum formula, we calculate the mean first return time ${\mathbb{E}}[T_{\rm FR}]$. Surprisingly, ${\mathbb{E}}[T_{\rm FR}]$ coincides with the result obtained from Kac’s lemma that applies to simple random walks (RWs). We also calculate the variance ${\rm Var}(T_{\rm FR})$, which accounts for the variability of first return times between different NBW trajectories. We apply this formalism to Erdős-Rényi networks, random regular graphs and configuration model networks with exponential and power-law degree distributions and obtain closed-form expressions for $P(T_{\rm FR}>t)$ as well as its mean and variance. These results provide useful insight on the advantages of NBWs over simple RWs in network exploration, sampling and search processes.
@@ 1 Introduction
Random walk (RW) models [1] provide useful tools for the analysis of dynamical processes on random networks [2, 3, 5, 4, 6, 7]. Here we focus on the case of undirected networks. Starting at time $t=0$ from a random initial node $i$, at each time step $t\geq 1$ an RW (also referred to as a simple RW) hops randomly to one of the neighbors of its current node. In some of the time steps the RW visits nodes that have not been visited before, while in other time steps it revisits nodes that have already been visited at an earlier time. The mean number $\langle S\rangle_{t}$ of distinct nodes visited by an RW on a random network up to time $t$ was studied in Ref. [8]. It was found that in the infinite network limit, in which random networks exhibit a tree structure, $\langle S\rangle_{t}\simeq rt$, where the coefficient $r<1$ depends on the network topology. In this case, the revisits are due to backtracking steps in which the RW hops back to the previous node and subsequent retroceding steps in which it keeps hopping backwards along its own path [9].
@@
In order to perform systematic studies of random walks on random networks, it is useful to focus on configuration model networks. The configuration model is an ensemble of uncorrelated random networks consisting of $N$ nodes, whose degree sequences are drawn from a given degree distribution $P(k)$. The admissible degrees are often restricted to a finite range $k_{\rm min}\leq k\leq k_{\rm max}$, where $k_{\rm min}$ is the minimal degree and $k_{\rm max}$ is the maximal degree, such that for any value of $k$ outside this range $P(k)=0$. The mean degree $\langle K\rangle$ is denoted by $c$. To ensure that an RW starting from any initial node $i$ will be able to reach any other node $j$, we focus on the case in which the whole network consists of a single connected component. Using the terminology of percolation theory, these are networks in which the giant component encompasses the whole network. In the large network limit, a sufficient condition for a configuration model network to consist of a single connected component is that $k_{\rm min}\geq 3$ [10, 6]. In fact, a weaker condition of $k_{\rm min}\geq 2$ is sufficient in the large network limit as long as a finite fraction of the nodes satisfy $k\geq 3$ [11]. We thus avoid isolated nodes of degree $k=0$ and leaf nodes of degree $k=1$, which may form isolated tree structures.
@@
The first return (FR) time $T_{\rm FR}$ of an RW is the first time at which it returns to the initial node $i$ [12]. The first return time varies between different instances of the random walk trajectory and its properties can be captured by a suitable distribution. The distribution of first return times may depend on the specific realization of the random network and on the choice of the initial node $i$. The distribution of first return times from a random node to itself in a given ensemble of random networks is denoted by $P(T_{\rm FR}=t)$. A classical result regarding first return times is Kac’s lemma, which states that the mean first return time of an RW from a given node $i$ to itself is given by [13, 14, 15]
@@
where $\langle X\rangle$ is the average of the random variable $X$ over the degree distribution $P(k)$.
@@
An important variant of the RW model is the non-backtracking random walk (NBW), in which the move backwards to the previous node is excluded [27]. Since backtracking steps are excluded, in the infinite network limit in which the network exhibits a tree structure, an NBW never revisits a previously visited node. In particular, it never returns to the initial node. In a finite network, the first return process of NBWs takes place only via non-retroceding trajectories, which rely on the existence of cycles.
@@
NBWs are important for several reasons, which are summarised below. They provide a more efficient way to explore and analyze complex networks, compared to standard random walks. This is due to the fact that by avoiding backtracking steps they can cover more of the network in less time [28]. NBWs are useful for identifying community structures within networks [29]. In certain types of networks, NBWs can help mitigate localization effects that might trap standard random walks in specific regions of the network.
The paper is organized as follows. In Sec. 2 we present the configuration model networks, their construction and essential properties. In Sec. 3 we present the non-backtracking random walk. In Sec. 4 we derive formulae for the distribution $P(T_{\rm FR}>t)$ of first return times of NBWs on configuration model networks and for its mean ${\mathbb{E}}[T_{\rm FR}]$ and variance ${\rm Var}(T_{\rm FR})$. In Sec. 5 we apply these results to RRGs, Erdős-Rényi (ER) networks and to configuration model networks with exponential and power-law distributions. The results are discussed in Sec. 6 and summarised in Sec. 7.
2 Configuration model networks
The configuration model is an ensemble of uncorrelated random networks whose degree sequences are drawn from a given degree distribution $P(k)$ [31, 33, 34, 32, 35]. These networks are simple graphs in the sense that each pair of nodes is connected by at most a single edge and there are no self-loops. The first moment (mean degree) and the second moment of $P(k)$ are denoted by $\langle K^{n}\rangle$, where $n=1$ and $2$, respectively, while the variance is given by $Var(K)=\langle K^{2}\rangle-\langle K\rangle^{2}$. The support of the degree distribution of random networks is often bounded from below by $k_{\rm min}\geq 1$ such that $P(k)=0$ for $0\leq k\leq k_{\rm min}-1$, with non-zero values of $P(k)$ only for $k\geq k_{\rm min}$. For example, the commonly used choice of $k_{\rm min}=1$ eliminates the possibility of isolated nodes in the network. Choosing $k_{\rm min}=2$ also eliminates the leaf nodes. One may also control the upper bound by imposing $k\leq k_{\rm max}$. This may be important in the case of finite networks with heavy-tail degree distributions such as power-law distributions. The configuration model network ensemble is a maximum entropy ensemble under the condition that the degree distribution $P(k)$ is imposed [33, 34, 32]. Here we focus on the case of undirected networks.
To generate a network instance drawn from an ensemble of configuration model networks of $N$ nodes, with a given degree distribution $P(k)$, one draws the degrees of the $N$ nodes independently from $P(k)$. This gives rise to a degree sequence of the form $k_{1},k_{2},\dots,k_{N}$. For the discussion below it is convenient to list the degree sequence in a decreasing order of the form $k_{1}\geq k_{2}\geq\dots\geq k_{N}$. It turns out that not every possible degree sequence is graphical, namely admissible as a degree sequence of a network. Therefore, before trying to construct a network with a given degree sequence, one should first confirm the graphicality of the degree sequence. To be graphical, a degree sequence must satisfy two conditions. The first condition is that the sum of the degrees is an even number, namely $\sum_{i}k_{i}=2L$, where $L$ is an integer that represents the number of edges in the network. The second condition is expressed by the Erdős-Gallai theorem, which states that an ordered sequence of the form $k_{1}\geq k_{2}\geq\dots\geq k_{N}$ that satisfies the first condition is graphical if and only if the condition
$\sum_{i=1}^{n}k_{i}\leq n(n-1)+\sum_{i=n+1}^{N}\min(k_{i},n)$ (4)
holds for all values of $n$ in the range $1\leq n\leq N-1$ [36, 37].
To construct a network instance consisting of $N$ nodes with a given degree sequence $k_{1},k_{2},\dots,k_{N}$ (where $\sum_{i=1}^{N}k_{i}=2L$ and $L$ is the total number of undirected edges), we create a multiset of $2L$ stubs which includes $k_{i}$ stubs for each node $i$. Pairs of stubs are then selected randomly and connected to each other to form edges between the corresponding nodes. To illustrate the process we represent the stubs by $2L$ balls, where the $k_{i}$ balls associated with node $i$ are marked by $i$. We then choose a random arrangement of the $2L$ balls in an array of $L$ cells, such that each cell includes exactly two balls. In practice, a random arrangement of balls into cells can be obtained by generating a random permutation of the $2L$ balls and grouping them sequentially into $L$ pairs, making the construction straightforward to implement. A cell containing balls $i$ and $j$ represents an edge between nodes $i$ and $j$. The representation in terms of balls and cells is particularly convenient for implementation on the computer, since a single random permutation of the $2L$ balls produces a uniformly random pairing of stubs, from which the network can be constructed directly.
The network obtained from the procedure described above is a multigraph with the given degree sequence, which may include self-loops (edges connecting a node to itself) or multiple edges (two or more edges connecting the same pair of nodes). To eliminate the self-loops and multiple edges, we apply an edge switching process, which yields a simple graph while preserving the degree sequence. In this process, as long as the network has not yet become a simple graph, at each time step we select randomly one of the self-loops $(i,i)$ or one of the multiple edges $(i,j)$. In case that a self-loop $(i,i)$ was selected, we select a random edge $(i^{\prime},j^{\prime})$ and swap the two edges into $(i,i^{\prime})$ and $(i,j^{\prime})$. Similarly, in case that a multiple edge $(i,j)$ was selected, we select a random edge $(i^{\prime},j^{\prime})$ and swap the two edges into $(i,i^{\prime})$ and $(j,j^{\prime})$. In both cases, we complete the move only after we make sure that the swapping does not create a new self-loop or a new multiple edge. This random edge-switching process continues until no self-loops or multiple edges remain. The procedure described above provides the random simple graph ensemble used in the simulations.
The elimination of multiple edges may introduce some degree-degree correlations in the resulting simple graph. To keep these degree-degree correlations negligible, the degree distribution must exhibit a structural cutoff such that the expected number of nodes of degree $k>\sqrt{Nc}$ is $o(1)$ [38, 39]. In the case of fat-tailed degree distributions such as the power-law degree distribution, one needs to impose an upper cutoff $k_{\rm max}<\sqrt{Nc}$. The degree distributions of all the network models considered here satisfy the above conditions, so degree-degree correlations are negligible.
Some commonly studied configuration model networks can be described in terms of single parameter families of degree distributions. These include the RRG, the ER network and configuration model networks with exponential and power-law degree distributions. A particularly convenient choice of the parameter is the mean degree $c=\langle K\rangle$. In this case, the degree distribution can be expressed by $P(k)=P_{c}(k)$, such that small values of $c$ correspond to the dilute network limit while large values of $c$ correspond to the dense network limit.
Configuration model networks in which the lower bound of the degree distribution satisfies $k_{\rm min}=0$ or $1$, may exhibit a percolation transition at some value $c_{0}$ of the mean degree, referred to as the percolation threshold. Below the transition the network consists of finite tree components, while above the transition a giant component emerges. The percolation transition is a second order phase transition, whose order parameter is the fraction $g$ of nodes that reside on the giant component. Below the transition, where $c<c_{0}$, the order parameter is $g=0$, while for $c>c_{0}$ the fraction $g=g(c)$ of nodes that reside on the giant component gradually increases. The giant component of a configuration model network consists of a $2$-core which is decorated by tree branches [40]. The $2$-core is a connected component, such that each node on the $2$-core has links to at least two other nodes that reside on the $2$-core. The nodes that reside on the tree branches have the property that their deletion would break the giant component into two or more components. Such nodes are referred to as articulation points [41, 42]. Similarly, the deletion of an edge that resides on one of the tree branches would break the giant component into two components. Such edges are referred to as bredges [43].
In this paper we focus on the case in which the whole network consists of a single connected component, for which $g=1$. Below we discuss the conditions for $g=1$ in RRGs, ER networks and configuration model networks with exponential and power-law distributions.
Consider an RRG that consists of $N$ nodes of degree $c$ (where $Nc$ is even). For $c=1$ the nodes form dimers. For $c=2$ the network, which is referred to as a $2$-random regular graph ($2$-RRG), consists of closed loops or cycles. In the large network limit, the expected number of cycles is $N_{C}\simeq\frac{1}{2}\ln N$ and the cumulative distribution of cycle lengths is given by $P(L\leq\ell)\simeq\ln\ell/\ln N$, where $\ell\leq N$ [44]. Here we focus on RRGs with $c\geq 3$, which in the large network limit consist of a single connected component [45, 46].
In the case of ER networks, in the large network limit there is a phase transition at $c_{1}=\ln N$, where $g\rightarrow 1$ [10, 6]. Above this point, the giant component encompasses the whole network, linking all the nodes into a single connected component.
In general, a sufficient condition for a configuration model network with degree distribution $P(k)$ to consist of a single connected component in the large network limit $N\rightarrow\infty$ is $k_{\rm min}\geq 3$ [47]. In fact, a weaker condition of $k_{\rm min}\geq 2$ is also sufficient, as long as a finite fraction of the nodes in the network are of degrees $k\geq 3$ [11]. In the analysis presented below of NBWs on configuration model networks with exponential and power-law distributions, we chose networks of size $N=1000$ that satisfy $k_{\rm min}\geq 3$. We checked each network instance to confirm that it consists of a single connected component.
In a finite configuration model network, there is a non-zero probability that the network will consist of a single connected component even if it includes some nodes of degree $k=1$. It was recently shown [48] that as the network size $N$ is increased, it may still consist of a single connected component with high probability as long as the number $n_{1}$ of nodes of degree $k=1$ grows more slowly than $\sqrt{N}$. However, in networks that include leaf nodes of degree $k=1$, NBWs that enter these nodes will get stuck. Therefore, in the study of NBWs it is important not only to ensure that the network consists of a single connected component, but also that this component does not include any leaf nodes. This implies that the $2$-core of the network (namely the largest subgraph in which all the nodes are of degree $k\geq 2$) encompasses the whole network. It also implies that the network does not include any articulation points [41, 42] or bredges [43].
3 Non-backtracking random walks
@@
NBWs are RWs for which the move backwards to the previous node is excluded. They belong to the class of second-order random walks, in which the transition probabilities depend not only on the current node but also on the previous node [27, 30]. This introduces memory into the process, which makes it no longer Markovian in the traditional sense. The challenge is to analyze such processes using methods that are typically applied to Markov chains, which rely on the memoryless property. Recently, Fasino et al. introduced a mapping of second order random walks into first order processes on a larger state space, referred to as the pullback process [30]. Instead of viewing the random walk as taking place between the nodes of the original graph, the pullback process considers a random walk on the directed-line graph associated with the original graph. Using this method they showed that the mean first return time ${\mathbb{E}}[T_{\rm FR}]$ of any second order random walk (including NBWs) on undirected networks satisfies Eq. (1), thus extending the validity of Kac’s lemma to second order random walks on undirected networks [30]. Note that Kac’s lemma deals with the mean first return time and has no implications on the overall shape of the distribution and its higher order moments.
@@
NBWs exhibit faster mixing times than standard random walks, meaning they converge to their stationary distribution more quickly [49, 50]. They thus inspire the design of more efficient algorithms for various graph-based problems, including link prediction and node centrality measures [51]. The non-backtracking (Hashimoto) matrix $B$ associated with these walks has spectral properties that can reveal important information about the network structure, often more clearly than traditional adjacency matrices [53, 52, 15]. Actually, the mixing time is inversely proportional to the spectral gap of the matrix $B$ [14]
@@
For the special case of RRGs it was shown that the mixing time of NBWs (and RWs) scales like $t_{\rm mix}\propto\ln N$ [28, 54]. This result was later generalized to a broader class of configuration model networks with $k_{\rm min}\geq 3$ [55]. This result sits well with the fact that both the mean distance [56, 57] and the diameter [61, 58, 59, 60, 62] of RRGs are proportional to $\ln N$. It implies that an NBW starting from a random initial node $i$ at time $t=0$ may reach any other node in the network within $\ln N$ time steps. Moreover, using the shell structure around the initial node $i$ as a spherical coordinate system, the radial component of the location of each node is given by its distance from $i$. Since RRGs are locally tree-like at distances in the range $\ell\ll\ln N$ [63], an NBW starting from $i$ essentially moves deterministically to the next shell away from $i$ as far as the tree-like structure persists. This is unlike the case of RWs which behave like biased random walks along the radial axis, moving outwards with probability 11/c1-1/c and inwards with probability $1/c$ [9, 57].
@@ 4 The distribution of first return times
In the first three time steps the NBW crosses new ‘directed’ edges with probability $1$, which implies that $\Delta L_{t}=1$ for $t=0,1$ and $2$. For $t\geq 3$ we use a mean-field approach, which essentially assumes that the ‘directed’ edges that have already been crossed and those that have not yet been crossed are distributed uniformly in the network and can thus be visited with equal probability at any time step. This approach applies under the condition that the network consists of a single connected component. A further condition is that the network will not be dominated by linear chains consisting of nodes of degree $k=2$, namely that $P(k=2)$ will be sufficiently small (networks that do not satisfy this condition are referred to as almost $2$-RRGs [11]).
In configuration model networks that consist of a single connected component with $k_{\rm min}\geq 3$ and no leaf nodes, the mixing time scales like $t_{\rm mix}\propto\ln N$ [54, 55], while the mean first return time scales like ${\mathbb{E}}[T_{\rm FR}]\propto N$. Thus, for sufficiently large networks $t_{\rm mix}\ll\mathbb{E}[T_{\rm FR}]$. This separation of time scales implies that apart from the very early stages of the first return trajectories, NBWs sample the ‘directed’ edges in a uniform fashion. Under these conditions, the probability that at time $t+1$ the NBW will cross a ‘directed’ edge which has been crossed before is equal to the fraction of ‘directed’ edges that have already been crossed. This fraction is given by $(\langle L\rangle_{t}-2)/(Nc-2)$, where the subtraction of $2$ from the numerator and the denominator accounts for the fact that the ‘directed’ edges crossed at times t1t-1 and $t$ cannot be crossed again at time $t+1$. This implies that the probability $\Delta L_{t}$ is given by
$\Delta L_{t}=1-\frac{\langle L\rangle_{t}-2}{Nc-2}.$ (7)
To simplify the analysis, we reduce Eq. (7) to the form
@@
The reduction from Eq. (7) to Eq. (8) relies on the assumption that the network is both sufficiently large and sufficiently dense, such that the product $Nc$ satisfies $Nc\gg 2$. In addition, this reduction becomes accurate when the expected number of ‘directed’ edges $\langle L\rangle_{t}$ which have already been visited by the NBW satisfies $\langle L\rangle_{t}\gg 1$. This condition is indeed satisfied for sufficiently long times. Since for $t\ll Nc$ the number of distinct ‘directed’ edges visited by the NBW satisfies $\langle L\rangle_{t}\simeq t$, the condition $\langle L\rangle_{t}\gg 1$ can be replaced by $t\gg 1$.
Inserting $\Delta L_{t}$ from Eq. (6) into Eq. (8), we obtain the recursion equation
$\langle L\rangle_{t+1}=\langle L\rangle_{t}\left(1-\frac{1}{Nc}\right)+1.$ (9)
Solving Eq. (9), we obtain
@@
While Eq. (8) is valid to a good approximation for $t\geq 3$, it becomes precise above the mixing time, where the random walker samples ‘directed’ edges in a uniform fashion. Thus, apart from the first few steps, Eq. (10) can be approximated by
@@
The probability that an NBW will not visit a specific random node of degree $k$ for the first time up to time $t$, can be expressed by $\left(1-\frac{\langle L\rangle_{t-1}}{Nc}\right)^{k}$. This is due to the fact that in order to visit a node of degree $k$ the NBW must enter via one of the $k$ edges connected to $i$. Since an NBW quickly loses memory of its initial node, the probability of not returning to an initial node of degree $k$ up to time $t$ is the same as the probability not to visit any node of the same degree up to time $t$. Therefore, the tail distribution of first return times, under the condition that the initial node $i$ is of degree $k$, is given by
@@
and the second moment is given by
@@
This result coincides with Kac’s lemma, which is obtained from general properties of discrete stochastic processes [13, 30]. Eq. (23) also implies that conditioning on initial nodes of a given degree $k$, the mean first return time is given by
@@
This result goes beyond the generalization of Kac’s lemma for second-order random walks [30] and is valid for the specific case of the NBW.
@@
${\rm Var}(T_{\rm FR})\simeq N^{2}c^{2}\left(2\bigg\langle\frac{1}{K^{2}}\bigg\rangle-\bigg\langle\frac{1}{K}\bigg\rangle^{2}\right).$ (28)